OpenStudy (anonymous):

4. If A= (x+1)/(x-1) and B= (x-1)/(x+1) then A – B = 4. If A= (x+1)/(x-1) and B= (x-1)/(x+1) then A – B = @Mathematics

OpenStudy (anonymous):

(x+1)/(x-1) - (x-1)/(x+1) (x+1)-(x-1)/x^2-1 =2/x^2-1

OpenStudy (anonymous):

2 in terms of x

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

I too got same answer but in answer key it's given 4x/x^2-1

OpenStudy (anonymous):

oh sorry thats wrong

OpenStudy (anonymous):

I got it. I made a mistake in sign. Aditi it's - of (x-1)^2

OpenStudy (anonymous):

Yep , I got it =) sorry for the wrong Answer

OpenStudy (anonymous):

let's see: x+1/x-1 - x-1/x+1 ={( x^2 + 2x + 1) - (x^2 - 2x+1)} / x^2 - 1 = {2x - (-2x)}/x^2-1 = 4x / x^2-1

OpenStudy (anonymous):

no problem Aditi. Thanks to both of you..

OpenStudy (anonymous):

U're welcome..